# N-Agent Ad Hoc Teamwork

### v2

@@ A.4.1 Validating the Existence of Coordination Conventions
Tables 12, 13, 14, 15, and 16 display the full cross-play results for all MARL algorithms on all tasks. Note that the tables are reflected across the diagonal axis for viewing ease. The values on the off-diagonal (i.e., where the two algorithms are not the same) are the cross-play score, while the values shown on the diagonal are self-play scores, computed as described in App. A.3.3. The cross-play and self-play scores displayed are means and standard errors. The rightmost column reflects the row average and standard deviation of the cross-play (xp) scores corresponding to the test set of VDN, QMIX, IQL, IPPO, and MAPPO, excluding the self-play score. Thus, the rightmost column reflects how well on average the row algorithm can generalize to unfamiliar teams.
@@
Figure 12: Cross-play table for the mpe-pp task.
Figure 13: Cross-play table for the 5v6 task.
Figure 14: Cross-play table for the 8m task.
Figure 15: Cross-play table for the 3s5z task.
Figure 16: Cross-play table for the 10v11 task.
A.5 Cross-play performance as a function of the number of agents.
To provide more insight on how a non-NAHT team performance changes as the number of unseen agents increases, Figure 17 plots QMIX and IPPO’s test scores on 5v6 when paired with various methods, as a function of the number of unseen agents. Note that $k=0$ corresponds to teams entirely composed of the evaluated agents, while $k=5$ corresponds to teams entirely composed of the unseen teammates, and are actually self-play scores, whereas the other points are cross-play scores. These points appear on the plot to contextualize the performance of the mixed teams.
QMIX and IPPO have similar returns in self-play, but comparing the two figures, we observe a difference in the rate at which the cross-play returns decline. In particular, QMIX’s performance declines most around $k=3$ across a variety of methods, while IPPO’s performance is minimized at $k=2$.
Figure 17: Mean and std. error of QMIX (left) and IPPO (right) test returns on 5v6, in the presence of uncontrolled agents generated by various MARL learning algorithms, as the number of unseen agents increases from $k=0$ to a maximum of $k=5$. The unseen agents are generated by the second method in the legend name.
@@ A.6 Computing Infrastructure
All value-based algorithms (e.g. QMIX, VDN, IQL) were run without parallelizing environments, while policy gradient algorithms were run with parallelized environments. All methods were trained for 20M steps on all tasks. Each run took between 12-48 hours of compute, and used less than 2gB of GPU memory. Runs were parallelized to use computational resources most efficiently. The servers used in our experiments ran Ubuntu 18.04 with the following configurations:

### v3

@@ A.5.2 Validating the Existence of Coordination Conventions
Table 4: Cross-play results for the mpe-pp task.
Table 5: Cross-play results for the 5v6 task.
Table 6: Cross-play results for the 8m task.
Table 7: Cross-play results for the 3s5z task.
Table 8: Cross-play results for the 10v11 task.
Tables 4, 5, 6, 7, and 8 display the full cross-play results for all MARL algorithms and POAM-AHT, on all tasks. Note that the tables are reflected across the diagonal axis for viewing ease. The values on the off-diagonal (i.e., where the two algorithms are not the same) are the cross-play score, while the values shown on the diagonal are self-play scores, computed as described in App. A.3.3. The cross-play and self-play scores displayed are means and 95% confidence intervals. The rightmost column reflects the row average and 95% confidence interval of the cross-play (XP) scores corresponding to the test set of VDN, QMIX, IQL, IPPO, and MAPPO, excluding the self-play score. Thus, the rightmost column reflects how well on average the row MARL/AHT algorithm can generalize to the test set used throughout this paper.
@@
A.5.3 Generalization to Unseen Teammate Types
As discussed and experimentally verified in Section A.5.2, diverse coordination conventions in StarCraft and MPE Predator Prey can be generated by (1) running the same MARL algorithm with various random seeds, or (2) running various MARL algorithms. In Section 6.3, the out-of-distribution (OOD) teammates considered were generated by running MARL algorithms with different random seeds than those used to generate train-time teammates—in other words, generating diverse and unseen teammates using the first procedure specified above.
Figure 12: Returns achieved by POAM and IPPO-NAHT, when trained on MAPPO, QMIX, and IQL, and tested on IPPO, VDN.
Here, we generate OOD teammates using the second procedure. More precisely, the five MARL algorithms used to generate teammates may be divided into train/test sets. Figure 12 shows the results of such an experiment, where POAM and IPPO-NAHT are trained with MAPPO, QMIX, and IQL teammates, and tested with agents from IPPO and VDN. Similar to the results presented in Section 6.3, we find that POAM generally outperforms IPPO-NAHT for the unseen teammate types for all tasks.
A.5.4 Modeling Controlled and Uncontrolled Agents
Figure 13: Loss of POAM’s encoder-decoder on the mpe-pp task, separated by uncontrolled (left) and controlled agents (right).
POAM’s encoder-decoder (ED) models both controlled and uncontrolled agents. Since the controlled agents are updated during the training process, the encoder-decoder must deal with a “moving target". Thus, in principle, the problem of modeling the controlled agents is more challenging than modeling uncontrolled agents.
Fig. 13 shows the probability of predicting the correct action for the uncontrolled agents and controlled agents separately, on the mpe-pp task. Note that the action probabilities shown in Fig. 5 (right) of the main paper would be the average of the two plots shown in Fig. 13.
For uncontrolled agents (Fig. 13, left), we observe that the accuracy of action predictions for the uncontrolled agents increases much more consistently as training goes on, and is higher than that for the controlled agents. As expected, the ED is able to model the uncontrolled agents more easily than the nonstationary controlled agents.
A.5.5 Performance as a Function of the Number of Controlled Agents
To provide more insight on how an NAHT team’s performance changes as the number of controlled agents increases, Figure 14 displays the mean test returns of POAM and POAM-AHT as a function of the number of controlled agents, where the evaluation returns are averaged across five types of uncontrolled teammates: QMIX, VDN, IQL, MAPPO, and IPPO. Note that the self-play returns of POAM and POAM-AHT (i.e. where the number of controlled agents is maximal) are shown as horizontal lines, while the performances at $N=0$ correspond to the averaged self-play returns of the five uncontrolled team types.
Recall that POAM-AHT was trained on the $N=1$ scenario only, while POAM is trained on the full NAHT setting. As expected, POAM outperforms POAM-AHT for all values of $N>1$, while for $N=1$, the methods perform similarly. POAM-AHT’s performance declines as the number of controlled agents increases, which likely occurs because the evaluation setting becomes further from the training setting as $N$ increases.
Figure 14: Comparing the performance of POAM versus POAM-AHT, as the number of controlled agents varies. POAM and POAM-AHT agents are evaluated with the following set of uncontrolled agents: QMIX, VDN, IQL, MAPPO, IPPO.
@@ A.6 Computing Infrastructure
All value-based algorithms (e.g. QMIX, VDN, IQL) were run without parallelizing environments, while policy gradient algorithms were run with parallelized environments. All methods were trained for 20M steps on all tasks. Each run took between 12-48 hours of compute, and used less than 2gB of GPU memory. Runs were parallelized to use computational resources efficiently. The servers used for our experiments ran Ubuntu 20.04 with the following configurations: